# NON-CLASSICAL GENERATING SETS FOR FUCHSIAN SCHOTTKY GROUPS

### v6

@@ Abstract.
The goal of this article is to initiate the study of estimates of the non-classical Schottky structure in the discrete subgroups of the projective special linear group over the real numbers degree $2$. In fact, in this paper, we have investigated the non-classical generating sets in the Fuchsian Schottky groups on the hyperbolic plane with boundary. A Schottky group is considered non-classical if the curves used in the Schottky construction are Jordan curves (except the Euclidean circles). More precisely, in this manuscript, we have provided a structure of the rank $2$ Fuchsian Schottky groups with non-classical generating sets by utilizing two suitable hyperbolic Möbius transformations on the upper-half plane model. In particular, we have derived two non-trivial examples of Fuchsian Schottky groups with non-classical generating sets in the upper-half plane with the circle at infinity as the boundary.
@@ 1. INTRODUCTION
In 1974, Marden ([5], [6]) introduced the concept of non-classical Schottky groups with a non-constructive proof. In 1975, Zarrow [12] claimed that he had found an example of a rank $2$ non-classical Schottky group, but later it was proved to be classical by Sato [8]. More precisely, the Schottky group constructed by Zarrow [12] was, in fact, a classical Schottky group, but on a different set of generators, and the demonstration of this was the main result in Sato’s paper [8]. After that, in 1991 Yamamoto [11] explicitly presented an example of the rank $2$ non-classical Schottky group in the Riemann sphere. Then, in 2009 Williams [10] also delivered an example of the rank 2 non-classical Schottky group. So, in the literature, the works related to the non-classical structure that have been studied in Schottky groups are all in Kleinian flavor. In this article, we aim to investigate the Schottky structure in the Fuchsian Schottky groups with non-classical generating sets in Fuchsian flavor. In [2], Button proved that all Fuchsian Schottky groups are classical Schottky groups, but not necessarily on the same set of generators. Furthermore, Button in [2] gave an example (with figure) of a Fuchsian group which is Schottky on a particular generating set, but non-classical on those generators. On the other hand, Marden ([5], [6]) also established that, all Fuchsian Schottky groups are classical. Therefore, in this direction, it is quite a difficult and fascinating problem to examine the following question.
@@
In [9], we have provided the structure of arbitrary finite rank classical Fuchsian Schottky groups in the hyperbolic plane with a boundary on the harmony of the real Schottky groups with two subsequent additional conditions:
@@
After that, we organized the construction of any finite rank generalized Fuchsian Schottky group in the upper-half plane with orientation preserving and orientation reversing isometries as side-pairing transformations. Furthermore, we have created interesting surfaces like finite Loss Ness monster and finite Jacob’s ladder from Fuchsian Schottky groups. The Fenchel-Nielsen coordinates in Tiemüller space for any arbitrary rank Fuchsian Schottky groups are also derived (see, [9] for more details).
@@
The preceding Figure $1$ supplies the structure of the rank $2$ classical Fuchsian Schottky group in the hyperbolic plane.
@@
It is well known that a Fuchsian group is a discrete subgroup of $PSL(2,\mathbb{R})$. On the other hand, a Schottky group is a special type of Kleinian group whereas the Kleinian groups are the discrete subgroups of $PSL(2,\mathbb{C})$ (see, [7] and [3] for details). In this paper, our goal is to set up the structure of the Fuchsian Schottky groups with non-classical generating sets by using the above two conditions $(a)$ and $(b)$ in the upper-half plane model with the circle at infinity as the boundary (see, Figure 2). In particular, in this article, after looking at the strategy performed by Yamamoto in [11], we explicitly present the structure of the rank $2$ Fuchsian Schottky groups with non-classical generating sets by imposing some new ideas in the technique operated by Yamamoto in the manuscript [11]. Firstly, for the rank $2$ Schottky group, Yamamoto [11] pointed out a suitable positive real number (viz., $\sqrt{2}$) in his used hyperbolic Möbius transformations (one of the generators out of two) so that the group became non-classical in the extended complex plane. But in this investigation, for rank $2$ Fuchsian Schottky groups we have observed that there doesn’t exist an analogous positive real number such that the groups become non-classical in the upper-half plane. Although the presence of negative real numbers is identified. For the lacuna of the existence of such a positive real number, in this study, we have taken the ‘$distance$’ (in the Euclidean sense) between the origin and the particular number (to specify the position of that number in the hyperbolic plane, i.e., the modulus values of the numbers, without loss of generality of the construction) in the used Möbius transformations (count as the rank of that group) to become these types of Schottky groups non-classical. Secondly, Yamamoto’s construction [11] was fabricated by two circles and two rectangles in the Riemann sphere where circles were occupied by the reflection of the imaginary axis and rectangles were built up with some kind of dilation (with rotation). On the other hand, here we have approached the construction of non-classical Fuchsian Schottky groups on the upper-half plane with four semi-circles centered on the circle at infinity, and all the four semi-circles are situated by the reflections of the upper imaginary axis. Thirdly, in Yamamoto’s paper [11], the centers of the circles are lying on the real axis whereas the midpoints of the rectangles are the same and it is the origin of the axis. However, the centers of all the semi-circles in our construction belong to $\mathbb{R}-[-1,+1]$. Further, the defining curves for Yamamoto’s non-classical Schottky group form the Jordan curves, whereas, in non-classical Fuchsian Schottky construction the semi-circles together with the diameters lying on the real axis create the Jordan loops look like the hollow half-moons (see, the red color curves in Figure $2$). In fact, in this manuscript, we have utilized these four half-moons as the defining curves to create the non-classical Fuchsian Schottky groups in the hyperbolic plane. Fourthly, Yamamoto [11] has applied the sufficiently small positive number in one generator out of two, but the nature of the construction of Fuchsian Schottky groups demands to use of the sufficiently small positive number in both the generators to become the group non-classical which we have provided in this work. These are the four new notions that we have imposed in this paper. In this way, we have developed the literature by explicitly setting up the hollow half-moons as Jordan curves in the Schottky structure for Fuchsian Schottky groups and produced two non-trivial examples of such groups with non-classical generating sets in the upper-half plane with boundary.
@@
In this paper, our aim is to prove the following Theorem 1 which indicates the structure of rank $2$ Fuchsian Schottky groups with generating sets that are non-classical on the hyperbolic plane. After that, as a consequence of Theorem 1, we deduce Corollary 1.0.1 and 1.0.2 that provide the two non-trivial examples of such groups in the upper-half plane with the circle at infinity as the boundary.
[…]
@@
[…]
@@
[…]
Now, we are going to state the main theorem of this paper.
Theorem 1..
Corollary 1.0.1.
Corollary 1.0.2.
In Section $4$, the elements ${h^{1}}_{(3,2)_{\kappa_{1}}}$, ${h^{2}}_{(4,1)_{\kappa_{1}}}$, ${h^{3}}_{(3,2)_{\kappa_{2}}}$, and ${h^{4}}_{(4,1)_{\kappa_{2}}}$ are described in detail.
@@
We take, for $\kappa\rightarrow 0^{+}$, $\Big{(}\frac{(\lambda+2)z+\{(\lambda+2)^{2}-(1-\kappa)^{2}\}}{z+(\lambda+2)}\Big{)}<(\lambda+2)$, $\forall$ $z$in $\mathbb{H}$ in the sense of the modulus values (i.e., the distance measurement) of the complex numbers in the upper-half plane, where $\mathbb{H}$ denotes the hyperbolic plane.
For example, let, $z$ $=$ $x+iy$, where, $x$in $\mathbb{R}$, $y$in $\mathbb{R}^{+}$.
@@
Outline of this paper. In Section $2$, we have exhibited the structure of the rank $2$ Fuchsian Schottky group with non-classical generating sets (denoted by $\Gamma^{N}_{S_{\kappa}}$) on the upper-half plane through five lemmas. Lemma 2.1 deals with the existence of a fundamental domain for the rank $2$ classical Fuchsian Schottky group which is surrounded by four semi-circles centered on the extended real numbers. In Lemma 2.2, under certain assumptions, we have computed the lengths of components of the domain of discontinuity for the group $\Gamma^{N}_{S_{\kappa}}$ that intersects the real axis. In Lemma 2.3, we have again estimated the lengths of components of the domain of discontinuity for the group $\Gamma^{N}_{S_{\kappa}}$ which intersects the upper vertical axis. Lemma 2.4 is related to the regions in Lemma 2.2 and Lemma 2.3 with the distance between the semi-circles in the set ${SC}^{*}_{\kappa}$, where ${SC}^{*}_{\kappa}=\{SC^{1}_{\kappa},SC^{2}_{\kappa},SC^{3}_{\kappa},...,SC^{\textbf{I}^{+}}_{\kappa}\}$, ($\textbf{I}^{+}$ denotes the positive natural number) is the complete list of the images of the semi-circles $SC_{1,\kappa},SC_{2,\kappa},SC_{3,\kappa}$, and $SC_{4,\kappa}$ applying by the group $\Gamma^{N}_{S_{\kappa}}$. In Lemma 2.5, we have displayed that the distance between the consecutive semi-circles in the set ${SC}^{*}_{\kappa}$ is less than $\frac{1}{5}$. Here, we have measured the distance along the real axis or the upper imaginary axis. After proving all these lemmas, we have established the main theorem of this article in Section $3$ (see, Theorem 1). Section $4$ deals with the two non-trivial examples of Fuchsian Schottky groups with non-classical generating sets in the upper-half plane model (see, Corollary 1.0.1 and 1.0.2).
2. THE KEY LEMMAS
@@
Suppose $F_{\Gamma_{S_{2}}}$ (see, the light green color region in Figure $3$) is the fundamental domain of $\Gamma_{S_{2}}$ encircled by four semi-circles $SC_{1},SC_{2},SC_{3}$, and $SC_{4}$, ($SC_{4}$ is paired with $SC_{1}$, whereas $SC_{3}$ twins with $SC_{2}$) containing four free edges. Note that, all the vertices (viz., $v_{1}$, $v_{2}$, $v_{3}$, $v_{4}$, $v_{5}$, $v_{6}$, $v_{7}$, and $v_{8}$) of the polygon that represents the Dirichlet fundamental region are neither proper nor improper (see, [1] for definitions of proper and improper vertices). Let, {}$\{\infty\}\in\Gamma_{S_{2}}$, as an ordinary point of the group. Obviously, $F_{\Gamma_{S_{2}}}$ produces an infinite area unbounded region in the upper-half plane. In [9], we have seen that the group $\Gamma_{S_{2}}$ is a discrete subgroup of $PSL(2,\mathbb{R})$ acting in the hyperbolic plane $\mathbb{H}^{2}$ =$=\{(x,y):x\in\mathbb{R},$ $y>0\}$. Let, ${SC}^{*}_{1}$, ${SC}^{*}_{2}$, ${SC}^{*}_{3}$, and ${SC}^{*}_{4}$ be the $4$ half disks (all are semi-circular in shape) in $\mathbb{H}^{2}\cup\{\infty\}$ centered on the extended real numbers with boundary $SC_{1},SC_{2}$, $SC_{3}$, and $SC_{4}$ respectively perpendicular to ${}\mathbb{R}\cup\{\infty\}$. Then the unbounded region ${F}^{*}_{\Gamma_{S_{2}}}$ of $\mathbb{H}^{2}$ is enclosed by four semi-circular disks ${SC}^{*}_{1}$, ${SC}^{*}_{2}$, ${SC}^{*}_{3}$, and ${SC}^{*}_{4}$ is a fundamental domain for $\Gamma_{S_{2}}$ in $\mathbb{H}^{2}\cup\{\infty\}$. Since ${h}_{s}$ is hyperbolic, we can consider a point ‘${\eta}$’ on the axis of the element ${h}_{s}$. So, there exists an element ${g}_{s}\in{\Gamma}_{{S_{2}}}$ such that ${g}_{s}({\eta})\in{F}^{*}_{\Gamma_{S_{2}}}$. Note that, in the hyperbolic plane the region ${F}^{*}_{\Gamma_{S_{2}}}$ is geodesically convex. Hence, one of the semi-circular half disks ${SC}^{*}_{1}$, ${SC}^{*}_{2}$, ${SC}^{*}_{3}$, and ${SC}^{*}_{4}$, say ${SC}^{*}_{1}$, separates the fixed points of $g_{s}h_{s}(g_{s})^{-1}$. Therefore, the boundary of the half disk ${SC}^{*}_{1}$, i.e., the semi-circular curve ${SC}_{1}$ will also separate these fixed points. So the required fundamental domain of $\Gamma_{S_{2}}$ is the zone surrounded by $(g_{s})^{-1}(SC_{1})$, $(g_{s})^{-1}(SC_{2})$, $(g_{s})^{-1}(SC_{3})$, and $(g_{s})^{-1}(SC_{4})$ which is basically the infinite area region $F_{{\Gamma}_{S_{2}}}$ in the upper-half plane model. This completes the proof. ∎
@@
Observe that, if two semi-circles in the set ${SC}^{*}_{\kappa}$ (say, ${SC}^{m_{1}}$ and ${SC}^{m_{2}}$, where $m_{1},m_{2}\in\textbf{I}^{+}$) are not separated by other semi-circles in the set ${SC}^{*}_{\kappa}$, these two semi-circles occupy the boundary of $F_{\Gamma^{N}_{S_{\kappa}}}$. Now, the semi-circles $SC^{j}_{\kappa}$ and $SC^{j+1}_{\kappa}$ together with any two semi-circles in the family of curves $\{q(SC_{i,{\kappa}}):q\in\Gamma^{N}_{S_{\kappa}},i=1,2,3,4\}$ surround a fundamental domain for the group $\Gamma^{N}_{S_{\kappa}}$. Suppose that $Y_{jk{\kappa}}=\{SC^{j}_{\kappa}\cap i^{k-1}\mathbb{R}^{+}\}$, where $k=1,2,3,$ and +=$\mathbb{R}^{+}=\{x\in\mathbb{R}:x>0\}$. In the hyperbolic plane, we set the largest distance between two consecutive semi-circles in the set ${SC}^{*}_{\kappa}$ as $Z_{j{\kappa}}=\max_{k}\{\lvert Y_{jk{\kappa}}-Y_{(j+1)k{\kappa}}\rvert\}$. Assume that $\Omega(\Gamma^{N}_{S_{\kappa}})$ is the domain of discontinuity of the group $\Gamma^{N}_{S_{\kappa}}$.
@@
The length of each component of $\{\Omega(\Gamma^{N}_{S_{\kappa}})\cap\mathbb{R}\}$ which meets the region made up of the union of intervals given as $[-(\lambda+3),-(\lambda+1)]\cup[(\lambda+1),(\lambda+3)]$ is less than the subsequent quantity
@@
where $\lambda=\lvert x\rvert>1$ and $\lvert x\rvert$ denotes the modulus value of the number $x$
@@
Let, $\{{h^{*}}_{(3,2)_{\kappa}}{h^{**}}_{(4,1)_{\kappa}}{h^{*}}_{(3,2)_{\kappa}}({h^{**}}_{(4,1)_{\kappa}})^{-1}\}(z)$ $=$ $D(z)$ (say).
@@
$+72x+20)+2\kappa(6x^{4}+16x^{3}+3x^{2}-16x-8)+\kappa^{2}(-2x^{4}-16x^{3}+x^{2}+48x+24)$
$+4\kappa^{3}(-x^{4}+x^{2}-8x-4)+\kappa^{4}(x^{4}-11x^{2}+8x+4)+6\kappa^{5}(x^{2})-2\kappa^{6}(x^{2})\textbf{\}}^{\frac{1}{2}}]$
Without loss of generality, we assume that, $09x^{4}+48x^{3}+91x^{2}+72x+20=$ 0.

### v7

# NON-CLASSICAL GENERATING SETS IN FUCHSIAN SCHOTTKY GROUPS

@@ Abstract.
The goal of this article is to initiate the study of estimates of the non-classical Schottky structure in the discrete subgroups of the projective special linear group over the real numbers degree $2$. In fact, in this paper, we have investigated the non-classical generating sets in the Fuchsian Schottky groups on the hyperbolic plane with boundary. A Schottky group is usually considered non-classical if the curves used in the Schottky construction are Jordan curves (except the Euclidean circles). More precisely, in this manuscript, we have provided a structure of the rank $2$ Fuchsian Schottky groups with non-classical generating sets by utilizing two suitable hyperbolic Möbius transformations on the upper-half plane model. In particular, we have derived two non-trivial examples of Fuchsian Schottky groups with non-classical generating sets in the upper-half plane with the circle at infinity as the boundary.
@@ 1. INTRODUCTION
In 1974, Marden ([5], [6]) introduced the concept of non-classical Schottky groups with a non-constructive proof. In particular, Marden [5] derived that the intersection of the closure of the classical Schottky space with the Schottky space is not the entire Schottky space, and hence there exist Schottky groups that are not classical. It is evident that a Schottky group is still classical if it is non-classical on a particular generating set. Although, it is also a very challenging task to construct a Schottky group with a non-classical generating set. In 1975, Zarrow [12] claimed that he had found an example of a rank $2$ non-classical Schottky group, but later it was proved to be classical by Sato [8]. More precisely, the Schottky group constructed by Zarrow [12] was, in fact, a classical Schottky group, but on a different set of generators, and the demonstration of this was the main result in Sato’s paper [8]. In 1991, Yamamoto [11] first presented an example of the rank $2$ non-classical Schottky group in the Riemann sphere. Then, in 2009 Williams [10] also delivered an example of the rank 2 non-classical Schottky group by applying the procedure suggested by Yamamoto in [11]. So, in the literature, the works related to the non-classical structure that have been studied in Schottky groups are all in Kleinian flavor. Indeed, in the Kleinian group theory, the non-classical structure of Schottky groups is known in the literature. However, in the Fuchsian group theory, the non-classical Schottky structure (more precisely, an explicit framework of a non-classical generating set) is still somewhat of an open problem. In $1998$, Button [2] proved that all Fuchsian Schottky groups are classical Schottky groups, but not necessarily on the same set of generators. Furthermore, Button in [2] furnished a brief example (with figure) of a Fuchsian group which is Schottky on a particular generating set, but non-classical on those generators. On the other hand, Marden ([5], [6]) also established that all Fuchsian Schottky groups are classical. Therefore, in this direction, it is quite a difficult and fascinating problem to examine the following question.
@@
Interestingly, it is known in the literature (more precisely, from Yamamoto’s work [11]) that, the defining curves for a rank $2$ non-classical Kleinian Schottky group are two circles and two rectangles lying in the extended complex plane. Now, in the Fuchsian group, in this paper, we introduce the non-classical Schottky structure with defining curves as four hollow half-moons, i.e., four semi-circles equipped with the diameters lying on the circle at infinity (indicated by the red color arcs in Figure $2$ in Section $2$). In particular, in the Fuchsian group theory, these four hollow half-moons are the Jordan curves that represent the non-classical Schottky structure in the upper-half plane. In essence, Zarrow [12] first provided the non-classical generating sets in the Kleinian Schottky group. In this article, we are going to initiate the non-classical generating sets in the Fuchsian Schottky group by establishing the following theorem.
Theorem 1..
In section $2$, the Möbius transformations ${h^{*}}_{(3,2)_{\kappa}}$ and ${h^{**}}_{(4,1)_{\kappa}}$ are described in detail.
As a consequence of Theorem 1, in the following, we deduce Corollary 1.0.1 and 1.0.2 that provide the two non-trivial examples of such groups in the upper-half plane with the circle at infinity as the boundary.
Corollary 1.0.1.
Corollary 1.0.2.
In Section $4$, the elements ${h^{1}}_{(3,2)_{\kappa_{1}}}$, ${h^{2}}_{(4,1)_{\kappa_{1}}}$, ${h^{3}}_{(3,2)_{\kappa_{2}}}$, and ${h^{4}}_{(4,1)_{\kappa_{2}}}$ are clarified in detail.
The paper is arranged as follows: In Section $2$, we have organized the set up for non-classical generating sets in Fuchsian Schottky groups in the hyperbolic plane. In Section $3$, we have exhibited the non-classical Schottky structure in the Fuchsian Schottky group of rank $2$ (denoted by $\Gamma^{N}_{S_{\kappa}}$) on the upper-half plane through five lemmas. Lemma 3.1 deals with the existence of a fundamental domain for the rank $2$ classical Fuchsian Schottky group which is surrounded by four semi-circles centered on the extended real numbers. In Lemma 3.2, under certain assumptions, we have computed the lengths of components of the domain of discontinuity for the group $\Gamma^{N}_{S_{\kappa}}$ that intersects the real axis. In Lemma 3.3, we have again estimated the lengths of components of the domain of discontinuity for the group $\Gamma^{N}_{S_{\kappa}}$ which intersects the upper vertical axis. Lemma 3.4 is related to the regions in Lemma 3.2 and Lemma 3.3 with the distance between the semi-circles in the set ${SC}^{*}_{\kappa}$, where ${SC}^{*}_{\kappa}=\{SC^{1}_{\kappa},SC^{2}_{\kappa},SC^{3}_{\kappa},...,SC^{\textbf{I}^{+}}_{\kappa}\}$, ($\textbf{I}^{+}$ denotes the positive natural number) is the complete list of the images of the semi-circles $SC_{1,\kappa},SC_{2,\kappa},SC_{3,\kappa}$, and $SC_{4,\kappa}$ applying by the group $\Gamma^{N}_{S_{\kappa}}$. In Lemma 3.5, we have displayed that the distance between the consecutive semi-circles in the set ${SC}^{*}_{\kappa}$ is less than $\frac{1}{5}$. Here, we have measured the distance along the real axis or the upper imaginary axis. After proving all these lemmas, we have established the main theorem of this article in Section $4$ (see, Theorem 1). Section $5$ deals with the two non-trivial examples of Fuchsian Schottky groups with non-classical generating sets in the upper-half plane model (see, Corollary 1.0.1 and 1.0.2).
2. TECHNICAL CORE FOR ORGANIZING NON-CLASSICAL STRUCTURE IN FUCHSIAN SCHOTTKY GROUPS
In [9], we have provided the structure of arbitrary finite rank classical Fuchsian Schottky groups in the hyperbolic plane with $\mathbb{R}$ $\cup$ $\{\infty\}$ as the boundary on the harmony of the real Schottky groups with two subsequent additional conditions (see, Figure $1$ in the following):
@@ 2. TECHNICAL CORE FOR ORGANIZING NON-CLASSICAL STRUCTURE IN FUCHSIAN SCHOTTKY GROUPS
Recall that, in this manuscript, we aim to initiate the non-classical structure (more precisely, non-classical generating sets) for the Fuchsian Schottky group of rank $2$ in the upper-half plane model.
@@
The preceding Figure $1$ supplies the structure of the rank $2$ classical Fuchsian Schottky group in the hyperbolic plane. Here, $SC_{1}$, $SC_{2}$, $SC_{3}$, and $SC_{4}$ indicate the semi-circles (with radius $r$ for $SC_{1}$, and $SC_{4}$; $R$ for $SC_{2}$, and $SC_{3}$) situating within the upper-half plane with centers lying on the boundary of the hyperbolic plane (i.e., circle at infinity). Here, $r$ and $R$ are the arbitrary positive real numbers. Apart from that, $h_{(4,1)}$, and $h_{(3,2)}$ represent the hyperbolic Möbius transformations that pair the semi-circles with each other.
@@
It is well known that a Fuchsian group is a discrete subgroup of $PSL(2,\mathbb{R})$ (see, [4]). On the other hand, a Schottky group is a special type of Kleinian group whereas the Kleinian groups are the discrete subgroups of $PSL(2,\mathbb{C})$ (see, [7] and [3] for details). In this paper, our goal is to set up the structure of the Fuchsian Schottky groups with non-classical generating sets by using the above two conditions $(a)$ and $(b)$ in the upper-half plane model with the circle at infinity as the boundary (see, Figure $2$ in the following). In particular, in this article, after looking at the strategy proposed by Yamamoto [11], we explicitly prepare the framework of the rank $2$ Fuchsian Schottky groups with non-classical generating sets by imposing some new ideas in the technique operated by Yamamoto in the manuscript [11]. Firstly, for the rank $2$ Schottky group, Yamamoto [11] pointed out a suitable positive real number (viz., $\sqrt{2}$) in his used hyperbolic Möbius transformations (one of the generators out of two) so that the group became non-classical in the extended complex plane. But in this investigation, for rank $2$ Fuchsian Schottky groups we have observed that there doesn’t exist an analogous positive real number such that the groups become non-classical in the upper-half plane. Although the presence of negative real numbers is identified. For the lacuna of the existence of such a positive real number, in this study, we have taken the ‘distancedistance’ (in the Euclidean sense) between the origin and the particular number (to specify the position of that number in the hyperbolic plane, i.e., the modulus values of the numbers, without loss of generality of the construction) in the used Möbius transformations (count as the rank of that group) to become these types of Schottky groups non-classical. Secondly, Yamamoto’s construction [11] was fabricated by two circles and two rectangles in the Riemann sphere where circles were occupied by the reflection of the imaginary axis and rectangles were built up with some kind of dilation (with rotation). On the other hand, here we have approached the construction of non-classical Fuchsian Schottky groups on the upper-half plane with four semi-circles centered on the circle at infinity, and all the four semi-circles are situated by the reflections of the upper imaginary axis. Thirdly, in Yamamoto’s paper [11], the centers of the circles are lying on the real axis whereas the midpoints of the rectangles are the same and it is the origin of the axis. However, the centers of all the semi-circles in our construction belong to $\mathbb{R}-[-1,+1]$. Further, the defining curves for Yamamoto’s non-classical Schottky group form the Jordan curves, whereas, in non-classical Fuchsian Schottky construction the semi-circles together with the diameters lying on the real axis create the Jordan loops look like the hollow half-moons (see, the red color curves in Figure $2$). In fact, in this manuscript, we have utilized these four hollow half-moons as the defining curves to organize the non-classical Schottky structure for Fuchsian Schottky groups in the hyperbolic plane. Fourthly, Yamamoto [11] has applied the sufficiently small positive number in one generator out of two, but the nature of the construction of Fuchsian Schottky groups demands to use of the sufficiently small positive number in both the generators to become the group non-classical which we have provided in this work. These are the four new notions that we have imposed in this paper. In this way, we have developed the literature by explicitly setting up the hollow half-moons as Jordan curves in the Schottky structure for Fuchsian Schottky groups and produced two non-trivial examples of such groups with non-classical generating sets in the upper-half plane with boundary.
$\text{ and }{h^{**}}_{(4,1)_{\kappa}}\text{ $(\kappa\rightarrow 0^{+})$ }:z$ $\longrightarrow$ $\frac{(\lambda+2)(1-\kappa)^{-1}z+(1-\kappa)\{(\lambda+2)^{2}(1-\kappa)^{-2}-1\}}{(1-\kappa)^{-1}z+(\lambda+2)(1-\kappa)^{-1}}$ (2)
@@
$\text{ and }SC_{4,\kappa}$ $=$ $\{\lvert z+(\lambda+2)\rvert=1-\kappa\}.$
@@
$\text{ and }{h^{**}}_{(4,1)_{\kappa}}(SC_{4,\kappa})$ $=$ $SC_{1,\kappa}.$
@@
We take, for $\kappa\rightarrow 0^{+}$, $\Big(\frac{(\lambda+2)z+\{(\lambda+2)^{2}-(1-\kappa)^{2}\}}{z+(\lambda+2)}\Big)<(\lambda+2)$, $\forall$ $z$ $\in$ $\mathbb{H}^{2}$ in the sense of the modulus values (i.e., the distance measurement) of the complex numbers in the upper-half plane, where $\mathbb{H}^{2}$ denotes the hyperbolic plane.
For example, let, $z$ == $x+iy$, where, $x$ $\in$ $\mathbb{R}$, $y$ $\in$ $\mathbb{R}^{+}$.
@@
3. PROOF OF THE LEMMAS
@@
Suppose $F_{\Gamma_{S_{2}}}$ (see, the light green color region in Figure $3$) is the fundamental domain of $\Gamma_{S_{2}}$ encircled by four semi-circles $SC_{1},SC_{2},SC_{3}$, and $SC_{4}$, ($SC_{4}$ is paired with $SC_{1}$, whereas $SC_{3}$ twins with $SC_{2}$) containing four free edges. Note that, all the vertices (viz., $v_{1}$, $v_{2}$, $v_{3}$, $v_{4}$, $v_{5}$, $v_{6}$, $v_{7}$, and $v_{8}$) of the polygon that represents the Dirichlet fundamental region are all improper vertices (see, [1] for the definition of improper vertex). Observe that, here, all vertices are infinite vertices of the overhead polygon. Let, $\{\infty\}\in\Gamma_{S_{2}}$, as an ordinary point of the group. Obviously, $F_{\Gamma_{S_{2}}}$ produces an infinite area unbounded region in the upper-half plane. In [9], we have seen that the group $\Gamma_{S_{2}}$ is a discrete subgroup of $PSL(2,\mathbb{R})$ acting in the hyperbolic plane $\mathbb{H}^{2}$ $\{(x,y):x\in\mathbb{R},$ y>0}y>0\}. Let, ${SC}^{*}_{1}$, ${SC}^{*}_{2}$, ${SC}^{*}_{3}$, and ${SC}^{*}_{4}$ be the $4$ half disks (all are semi-circular in shape) in $\mathbb{H}^{2}\cup\{\infty\}$ centered on the extended real numbers with boundary $SC_{1},SC_{2}$, $SC_{3}$, and $SC_{4}$ respectively perpendicular to $\mathbb{R}\cup\{\infty\}$. Then the unbounded region ${F}^{*}_{\Gamma_{S_{2}}}$ of $\mathbb{H}^{2}$ is enclosed by four semi-circular disks ${SC}^{*}_{1}$, ${SC}^{*}_{2}$, ${SC}^{*}_{3}$, and ${SC}^{*}_{4}$ is a fundamental domain for $\Gamma_{S_{2}}$ in $\mathbb{H}^{2}\cup\{\infty\}$. Since ${h}_{s}$ is hyperbolic, we can consider a point $\eta$}’ on the axis of the element ${h}_{s}$. So, there exists an element ${g}_{s}\in{\Gamma}_{{S_{2}}}$ such that ${g}_{s}({\eta})\in{F}^{*}_{\Gamma_{S_{2}}}$. Note that, in the hyperbolic plane the region ${F}^{*}_{\Gamma_{S_{2}}}$ is geodesically convex. Hence, one of the semi-circular half disks ${SC}^{*}_{1}$, ${SC}^{*}_{2}$, ${SC}^{*}_{3}$, and ${SC}^{*}_{4}$, say ${SC}^{*}_{1}$, separates the fixed points of $g_{s}h_{s}(g_{s})^{-1}$. Therefore, the boundary of the half disk ${SC}^{*}_{1}$, i.e., the semi-circular curve ${SC}_{1}$ will also separate these fixed points. So the required fundamental domain of $\Gamma_{S_{2}}$ is the zone surrounded by $(g_{s})^{-1}(SC_{1})$, $(g_{s})^{-1}(SC_{2})$, $(g_{s})^{-1}(SC_{3})$, and $(g_{s})^{-1}(SC_{4})$ which is basically the infinite area region $F_{{\Gamma}_{S_{2}}}$ in the upper-half plane model. This completes the proof. ∎
@@
Observe that, if two semi-circles in the set ${SC}^{*}_{\kappa}$ (say, ${SC}^{m_{1}}$ and ${SC}^{m_{2}}$, where $m_{1},m_{2}\in\textbf{I}^{+}$) are not separated by other semi-circles in the set ${SC}^{*}_{\kappa}$, these two semi-circles occupy the boundary of $F_{\Gamma^{N}_{S_{\kappa}}}$. Now, the semi-circles $SC^{j}_{\kappa}$ and $SC^{j+1}_{\kappa}$ together with any two semi-circles in the family of curves $\{q(SC_{i,{\kappa}}):q\in\Gamma^{N}_{S_{\kappa}},i=1,2,3,4\}$ surround a fundamental domain for the group $\Gamma^{N}_{S_{\kappa}}$. Suppose that $Y_{jk{\kappa}}=\{SC^{j}_{\kappa}\cap i^{k-1}\mathbb{R}^{+}\}$, where $k=1,2,3,$ and $\mathbb{R}^{+}=\{x\in\mathbb{R}:x>0\}$. In the hyperbolic plane, we set the largest distance between two consecutive semi-circles in the set ${SC}^{*}_{\kappa}$ as $Z_{j{\kappa}}=\max_{k}\{\lvert Y_{jk{\kappa}}-Y_{(j+1)k{\kappa}}\rvert\}$. Assume that $\Omega(\Gamma^{N}_{S_{\kappa}})$ is the domain of discontinuity of the group $\Gamma^{N}_{S_{\kappa}}$.
@@
The length of each component of $\{\Omega(\Gamma^{N}_{S_{\kappa}})\cap\mathbb{R}\}$ which meets the region made up of the union of intervals given as $(\lambda+3),-(\lambda+1)]\cup[(\lambda+1),(\lambda+3)]$ is less than the subsequent quantity
@@
Let, $\{{h^{*}}_{(3,2)_{\kappa}}{h^{**}}_{(4,1)_{\kappa}}{h^{*}}_{(3,2)_{\kappa}}({h^{**}}_{(4,1)_{\kappa}})^{-1}\}(z)$ == $D(z)$ (say).
@@
$72x+20)+2\kappa(6x^{4}+16x^{3}+3x^{2}-16x-8)+\kappa^{2}(-2x^{4}-16x^{3}+x^{2}+48x+24)$
$4\kappa^{3}(-x^{4}+x^{2}-8x-4)+\kappa^{4}(x^{4}-11x^{2}+8x+4)+6\kappa^{5}(x^{2})-2\kappa^{6}(x^{2})\textbf{\}}^{\frac{1}{2}}]$
Without loss of generality, we assume that, $9x^{4}+48x^{3}+91x^{2}+72x+20=0$.